OpenStudy (anonymous):

Define a variable , write an equation,and solve each problem. Then check your solution. 1. Six times a number is -96. find the number 2.One fourth of a number is -16.325 . What is the number.

OpenStudy (gw2011):

A variable is a term that can have any value and it is shown by a letter. 1) 6x=-96 This equation shows x as the variable x=-96/6=-16 The number is -16 2) (1/4)x=-16.325 x=(4)(-16.325)=-65.3 The number is -65.3

OpenStudy (gw2011):

To check your solution: 1) (6)(-16)=-96 2) (1/4)(-65.3)=-16.325